An University wishes to hire 4 faculty members for its newly established department. They have
to select from a group of 8 Male and 7 Female applicants, all of whom are equally qualified to
fill in the positions. If University selects 4 at random, what is the probability that ;
a) All four will be Men
b) All four will be Women
c) Atleast one will be a Women
d) Atleast one will be a Men
1
Expert's answer
2020-06-10T19:47:13-0400

The University wishes to hire 4 faculty members from a group of 8 Male and 7 Female applicants i.e. total 8 + 7 = 15 applicants.


This can be done in "15 \choose4" = 1365 number of ways, which is the number of total outcomes.


We have, the total number of ways of choosing a Men and b Women from a group of A Men and B Women is: "{A\choose a}{B\choose b}"


a) All four will be Men

Here, A = 8, B = 7, a = 4, b = 0


"\therefore" The number of ways of choosing 4 Men = "{8\choose 4}{7\choose 0}"= 70


"\therefore" The probability that all four will be Men = (no of favorable cases)/(no of total outcomes)


= 70/1365 = 0.0513 (rounded to 4 decimal places)


Answer: The probability that all four will be Men is 0.0513.


b) All four will be Women

Here, A = 8, B = 7, a = 0, b = 4


"\therefore" The number of ways of choosing 4 Women = "{8\choose 0}{7\choose 4}" = 35


"\therefore" The probability that all four will be Women = (no of favorable cases)/(no of total outcomes)


= 35/1365 = 0.0256 (rounded to 4 decimal places)


Answer: The probability that all four will be Women is 0.0256.


c) At least one will be a Women


The probability that at least one will be a Women

= 1 - The probability that there will be no Women [since, P(Ac) = 1 - P(A)]

= 1 - The probability that all four will be Men

= 1 - 0.0513 [by using answer a)]

= 0.9487


Answer: The probability that at least one will be a Women is 0.9487.


d) At least one will be a Men


The probability that at least one will be a Men

= 1 - The probability that there will be no Men [since, P(Ac) = 1 - P(A)]

= 1 - The probability that all four will be Women

= 1 - 0.0256 [by using answer b)]

= 0.9744


Answer: The probability that at least one will be a Men is 0.9744.